OpenStudy (anonymous):

I dont get this one, Helen Weller invested $14,000 in an account that pays 10% simple interest. How much additional money must be invested in an account that pays 13% simple interest so that the total interest is equal to the interest on the two investments at the rate of 11%?

OpenStudy (anonymous):

14(10)+x(13)=(14+x)11 140+13x=154+11x 2x=14 x=7 Invest $7000 in the thirteen percent investment.

OpenStudy (anonymous):

Thank you!!!!!!!